OpenStudy (gelow):

Estimate by first rounding each number to the nearest integer. 4 1/9 + 2 2/13 A. –7 B. –6 C. 6 D. 7

OpenStudy (gelow):

@AnswerMyQuestions

OpenStudy (gelow):

@Korey_Minter

OpenStudy (gelow):

@Catlover5925 @chosenmatt

OpenStudy (freckles):

1/9 is closer to zero then 1 2/13 is also closer to zero then 1 so 4 and 1/9 is closer to 4 or 5? and is is 2 and 2/13 closer to 2 or 3?

OpenStudy (gelow):

4 and 2?

OpenStudy (freckles):

yes so 4+2=?

OpenStudy (freckles):

9/2=4.5 so if we have had: \[4\frac{\text{ a number between 4.5 and 9 including 4.5 and 9 }}{9} \approx 5\] 13/2=6.5 so if we have had: \[2\frac{ \text{ a number between 6.5 and 13 including 6.5 and 13 }}{13} \approx 3\] for example: \[4 \frac{5}{9} \approx 5 \\ 2\frac{12}{13} \approx 3 \\ 4 \frac{4}{9} \approx 4 \\ 2\frac{5}{13} \approx 2\]